Reports 1-1 of 1 Clear search Modify search
VIS (General)
kouseki.miyo - 21:00 Tuesday 21 July 2020 (14794) Print this report
Preliminary result of new accelerometer

Takahashi, Miyo

Hardware preparation is posted in 14793

what we did
we prepared servo controller in the DGS system to operate the new accelerometer which is demonstrated in Mitaka with analog circuit, and measured the seismic noise by this new accelerometer.

  • We get eigen frequency of 0.24 Hz by tunning the weight of the proof mass and tune the position of LVDT sensor.
  • We install the LVDT driver (3130) and coil driver (7393) in IY0 rack.
  • We made a servo filter with lead-lag filter which has a pole and zero are 3Hz and 0.1Hz, respectively. UGF was 1.5Hz.
  • We confirmed that the accelerometer signal was consistent with the seismometer in IXV. (see attached figure or "/users/VIS/Acctest/ProtoType/compare_seismometer.xml"). In this comparison, we defined a calibration factor from count to um/sec is 0.00237 in K1:VIS-TEST_ACC_FB2ACC_GAIN so that the acc. signal should be same as the seismometer siggnal in the frequency band that has a coherence (almost 1).

what we do
We will tune the servo fileter to increase the band width, and prepare for long measurement.

  • we could not increase due to a weak actuation because the Modified LPCD was used in this test. So, we use HPCD (4827) tomorrow.
  • According to todays result, we cannot understand the what is the noise limiting the accelerometer signal below 0.1Hz. So we install long term (1 month?) to check whether the spectrum is limited by the seismic noise or sensor noise. We will prepare the cabling for this test tomorrow.
Search Help